# Lossy Beyond Diagonal Reconfigurable Intelligent Surfaces: Modeling and Optimization

### v1

@@ II-A.2 Forest-Connected Reconfigurable Admittance Network
In a forest-connected BD-RIS with $G$ uniform groups, each group $g$ follows a tree-connected architecture [10], which means each port is only directly connected to one other port via a tunable admittance component. This tree-connected architecture can take the form of either a tridiagonal RIS or an arrowhead RIS. Two 36-element forest-connected BD-RISs with a group size 3 in tridiagonal form and arrowhead form are illustrated in Fig. 1(b) and Fig. 1(c), respectively. In a tridiagonal BD-RIS, every pair of adjacent ports within the $\bar{M}$ ports is connected via a tunable admittance component. In an arrowhead BD-RIS, one designated port among the $\bar{M}$ ports connects to all other ports, while the remaining ports are connected only to the ground. Both tree-connected forms have the same circuit complexity. In this case, the circuit complexity of a forest-connected BD-RIS reduces to $(2\bar{M}-1)G$. The admittance matrix $\mathbf{Y}$ still satisfies (2) – (4). In addition, for the tridiagonal form, the following constraint is imposed:
@@
$Y_{m_{g},n_{g}}=0,~{}\forall m_{g},n_{g},~{}m_{g}\neq c_{g},n_{g}\neq c_{g},m_{g}\neq n_{g}.$ (6)
Based on the above illustrations, the D-RIS with a single-connected architecture where each port is connected to ground via its own admittance component without interacting with other ports is a special case of BD-RIS with both group and forest-connected architectures when $\bar{M}=1$. The simple architecture of D-RIS makes it possible to individually model the loss of each tunable component and reflect it directly to its diagonal scattering matrix $\bm{\Phi}$ [22]. However, this does not hold for BD-RIS, whose scattering matrix has coupled entries due to the interconnections. To tackle this issue, we use admittance parameters to facilitate the modeling of lossy BD-RIS. Applying admittance parameter allows us to establish a linear mapping between the tunable admittance components expressed by the circuit model and the admittance matrix $\mathbf{Y}$. This further benefits the analysis of the BD-RIS loss, which fundamentally comes from the resistance component in the circuit model. Consequently, modeling the scattering matrix can be converted into modeling each tunable admittance component.
@@ II-B Modeling of Lossy Tunable Admittance
In practice, each tunable admittance component is realized using a semiconductor device that can adjust its admittance. Given that the physical length of a reconfigurable admittance component is usually much smaller than the wavelength of the incident signal, the response of each tunable admittance component can be accurately characterized by an equivalent lumped circuit model, regardless of its geometry realization [27]. As illustrated in Fig. 1, each tunable admittance $Y_{m_{g},n_{g}}$ is modeled as a parallel resonant circuit  [^3] This circuit models a practical varactor diode, where the varactor is represented by a tunable capacitance $C$ in series with a resistor $R$ and an inductor $L_{2}$, along with a parallel inductor $L_{1}$. The resistor $R$ accounts for the parasitic resistance of the varactor, which includes the energy loss, while $L_{2}$ represents the parasitic inductance. From a practical design perspective, $L_{1}$ is incorporated to extend the tuning range of the circuit’s resonances. From an optimization perspective, $L_{1}$ is introduced to adjust the feasible susceptance values of each admittance component. [27]. This lumped circuit model, initially used in D-RIS to capture the relationship between reflection amplitude and phase shift for each reflecting element [22], is generalized in BD-RIS to model the losses of each admittance component. The admittance of this equivalent circuit is given by
$Y_{m_{g},n_{g}}(C_{m_{g},n_{g}})=\frac{1}{\jmath\omega L_{1}}+\frac{1}{\jmath\omega L_{2}+\frac{1}{\jmath\omega C_{m_{g},n_{g}}}+R},\forall m_{g},n_{g},$ (7)
@@
($\left(\Re\{Y_{m_{g},n_{g}}\}-\frac{1}{2R}\right)^{2}$ (9)
+($\quad+\left(\Im\{Y_{m_{g},n_{g}}\}-\left(-\frac{1}{\omega L_{1}}\right)\right)^{2}=\left(\frac{1}{2R}\right)^{2}.$
Clearly, the radius of this circle is $\frac{1}{2R}$, which depends solely on the resistance $R$, and the center of the circle is located at $\left(\frac{1}{2R},-\frac{1}{\omega L_{1}}\right)$, which depends jointly on the resistance $R$ and the inductance $L_{1}$. In this paper, we utilize the parameters from the data sheet of the practical varactor diode SMV2020-079 to model the equivalent circuit of each admittance component. Specifically, we set $L_{2}=0.7$ nH, $C_{m_{g},n_{g}}\in[0.35,3.20]$ pF [^4] The limited range of capacitance values, along with the presence of $L_{1}$ prevents the tunable admittances of interconnections from being completely ”turned off”, indicating that single-connected architecture is not a subset of the fully-/group-connected and tree-/forest-connected architectures in BD-RIS. and assume $L_{1}=6$ nH  [^5] In [22], $L_{1}$ is assumed to be 2.5 nH to achieve the full reflection tuning range for each reflecting element. However, this assumption is unnecessary in our BD-RIS model due to the nonlinear mapping from the admittance matrix to scattering matrix. The selection of the $L_{1}$ value is discussed in Section V.. As such, we represent the possible values of $Y_{m_{g},n_{g}}$ through the set $\mathcal{Y}$, defined as
$\mathcal{Y}=\{\alpha+\beta e^{\jmath\theta}\ |\ \theta\in[\theta_{\mathrm{min}},\theta_{\mathrm{max}}]\},$ (10)
@@
Remark 1: In most existing D-RIS [1, 2, 3] and most existing BD-RIS literature666[20] studies lossy connections between admittance components while all the admittance components themselves are assumed to be lossless. [5, 6, 17, 14, 8, 9, 10, 15, 13, 25, 18, 19, 21, 16, 11, 12, 26], the reconfigurable admittance network is assumed to be lossless, such that each tunable admittance is purely imaginary with $\Re\{Y_{m_{g},n_{g}}\}=0$ and the feasible range of its imaginary part lies unconstrained along the y-axis as illustrated in Fig. 2. This simple assumption fails to capture the coupling between the real and imaginary parts of $Y_{m_{g},n_{g}}$, which is important in practical realizations of BD-RIS and will have a non-negligible impact on the system performance as it will be illustrated in the following sections.
@@ III-A System Model and Problem Formulation
We start by considering an $M$-element BD-RIS-aided SISO wireless communication system. In this paper, we adopt the common assumptions from related literature [5, 6, 17, 14, 8, 9, 10, 15, 20, 21, 16, 11, 12, 26], which state that the antennas at the transmitter, RIS, and receiver are perfectly matched, have no mutual coupling and no structural scattering. The overall channel $h$ between the transmitter and receiver can thus be written as
@@
Next, we aim to maximize the received signal power for the considered BD-RIS-aided SISO system, taking into account the losses in BD-RIS. For simplicity of presentation, we focus on the group-connected architecture. The proposed algorithms can be readily adapted to the forest-connected architecture; see Section III-D. As discussed in Section II-A, it is convenient to capture the losses of BD-RIS with admittance parameter. By employing the relationship between the scattering matrix $\bm{\Phi}$ and the admittance matrix $\mathbf{Y}$ in (1), the mapping between $\mathbf{Y}$ and the tunable admittances $Y_{m_{g},n_{g}},\forall m_{g},n_{g}$ in (2) – (4), and the set of feasible values for each $Y_{m_{g},n_{g}}$ in (10), the power maximization problem can be formulated as follows:
@@
$\mathbf{Y}_{g}=\mathbf{Y}_{g}^{T},\forall g\in\mathcal{G},$ (12d)
@@
$Y_{m_{g},n_{g}}\in\mathcal{Y},\forall m_{g},n_{g}.$ (12f)
Problem (12) is challenging to solve because constraint (12b) introduces a nonlinear mapping between the admittance matrix and the scattering matrix, resulting from matrix inversion. Additionally, constraint (12f) is non-convex, which further complicates the optimization process. In the following subsections, we propose two solutions to maximize the received signal power with a lossy BD-RIS in the SISO systems. First, we propose an MM-ADMM algorithm in Section III-B to directly solve (12). In Section III-C, we further propose a low-complexity algorithm that approximately maximizes the received signal power without directly solving (12).
III-B Proposed MM-ADMM Algorithm
@@
Let $\bar{\mathbf{Y}}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$ be a matrix with entries $[\bar{\mathbf{Y}}_{g}]_{m,n}=Y_{m_{g},n_{g}}$, then (12e) can be expressed as a linear mapping from $\bar{\mathbf{Y}}_{g}$ to $\mathbf{Y}_{g},\forall g\in\mathcal{G}$. To facilitate the optimization, we further re-write this linear mapping into a vector form as $\mathsf{vec}(\mathbf{Y}_{g})=\mathbf{B}\mathsf{vec}(\bar{\mathbf{Y}}_{g})$, where matrix $\mathbf{B}\in\{0,\pm 1\}^{\bar{M}^{2}\times\bar{M}^{2}}$ maps $\mathsf{vec}(\bar{\mathbf{Y}}_{g})$ to $\mathsf{vec}(\mathbf{Y}_{g})$, $\forall g\in\mathcal{G}$, and is defined as
$[\mathbf{B}]_{\bar{M}(i-1)+j,l}=\begin{cases}1,&l=\bar{M}(k-1)+i~{}\text{and}~{}j=i;\\ -1,&l=\bar{M}(i-1)+j~{}\text{and}~{}j\neq i;\\ 0,&\text{otherwise},\end{cases}$ (16)
@@
$[\mathbf{P}]_{\bar{M}(i-1)+j,k}=$ (17)
$\ \begin{cases}1,&k=\frac{(2\bar{M}-j)(j-1)}{2}+i~{}\text{and}~{}1\leq j\leq i;\\ 1,&k=\frac{(2\bar{M}-i)(i-1)}{2}+j~{}\text{and}~{}i<j\leq\bar{M};\\ 0,&\text{otherwise},\end{cases}$
[…]
@@
$\mathrm{s.t.}~{}~{}~{}~{}~{}~{}$ $\bm{\Phi}_{g}=\left(Y_{0}\mathbf{I}_{\bar{M}}+\mathbf{Y}_{g}\right)^{-1}\left(Y_{0}\mathbf{I}_{\bar{M}}-\mathbf{Y}_{g}\right),\forall g\in\mathcal{G},$ (18b)
$\mathbf{Y}_{g}=\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}),\forall g\in\mathcal{G},$ (18c)
$[\bar{\mathbf{y}}_{g}]_{m^{\prime}}\in\mathcal{Y},\forall g\in\mathcal{G},\forall m^{\prime}\in\mathcal{M}^{\prime},$ (18d)
where =$\mathcal{M^{\prime}}=\{1,\ldots,\frac{\bar{M}(\bar{M}+1)}{2}\}$. To address this, we propose a double-loop iterative algorithm, referred to as the MM-ADMM algorithm. In the outer loop, we utilize the MM method [28] to linearize objective (18a). Then, in the inner loop, we employ the ADMM method [29] to effectively handle the complex constraints. The details of the proposed algorithm are described as follows.
III-B.1 MM Approximation
The main idea of the MM method for solving (18) is to replace the objective function with a locally tight lower-bound surrogate function, usually a concave one, such that the resulting problem is easier to tackle. For (18a), a simple and common approach for this approximation is to use first-order Taylor expansion of $\tilde{\bm{\phi}}^{H}\mathbf{F}\tilde{\bm{\phi}}$. Defining $g(\tilde{\bm{\phi}})=\tilde{\bm{\phi}}^{H}\mathbf{F}\tilde{\bm{\phi}}$, at iteration $t$, we have

### v2

@@ II-A.2 Forest-Connected Reconfigurable Admittance Network
In a forest-connected BD-RIS with $G$ uniform groups, each group $g$ follows a tree-connected architecture [nerinigraphtheory], which means each port is only directly connected to one other port via a tunable admittance component. This tree-connected architecture can take the form of either a tridiagonal RIS or an arrowhead RIS. Two 36-element forest-connected BD-RISs with a group size 3 in tridiagonal form and arrowhead form are illustrated in Fig. 1(b) and Fig. 1(c), respectively. In a tridiagonal BD-RIS, every pair of adjacent ports within the $\bar{M}$ ports is connected via a tunable admittance component. In an arrowhead BD-RIS, one designated port among the $\bar{M}$ ports connects to all other ports, while the remaining ports are connected only to the ground. Both tree-connected forms have the same circuit complexity. In this case, the circuit complexity of a forest-connected BD-RIS reduces to $(2\bar{M}-1)G$. The admittance matrix $\mathbf{Y}$ still satisfies (2) – (4). In addition, for the tridiagonal form, the following constraint is imposed:
@@
$Y_{m_{g},n_{g}}=0,~\forall m_{g},n_{g},~m_{g}\neq c_{g},~n_{g}\neq c_{g},~m_{g}\neq n_{g}.$ (6)
Based on the above illustrations, the D-RIS with a single-connected architecture where each port is connected to ground via its own admittance component without interacting with other ports is a special case of BD-RIS with both group and forest-connected architectures when $\bar{M}=1$. The simple architecture of D-RIS makes it possible to individually model the loss of each tunable component and reflect it directly to its diagonal scattering matrix $\bm{\Phi}$ [practicalRIS_Abeywickrama]. However, this does not hold for BD-RIS, whose scattering matrix has coupled entries due to the interconnections. To tackle this issue, we use admittance parameters to facilitate the modeling of lossy BD-RIS. Applying admittance parameter allows us to establish a linear mapping between the tunable admittance components expressed by the circuit model and the admittance matrix $\mathbf{Y}$. This further benefits the analysis of the BD-RIS loss, which fundamentally comes from the resistance component in the circuit model. Consequently, modeling the scattering matrix can be converted into modeling each tunable admittance component.
@@ II-B Modeling of Lossy Tunable Admittance
In practice, each tunable admittance component is realized using a semiconductor device that can adjust its admittance. Given that the physical length of a reconfigurable admittance component is usually much smaller than the wavelength of the incident signal, the response of each tunable admittance component can be accurately characterized by an equivalent lumped circuit model, regardless of its geometry realization [koziel2013surrogate]. As illustrated in Fig. 1, each tunable admittance $Y_{m_{g},n_{g}}$ is modeled as a parallel resonant circuit [^3] This circuit models a practical varactor diode, where the varactor is represented by a tunable capacitance $C$ in series with a resistor $R$ and an inductor $L_{2}$, along with a parallel inductor $L_{1}$. The resistor $R$ accounts for the parasitic resistance of the varactor, which includes the energy loss, while $L_{2}$ represents the parasitic inductance. From a practical design perspective, $L_{1}$ is incorporated to extend the tuning range of the circuit’s resonances. From an optimization perspective, $L_{1}$ is introduced to adjust the feasible susceptance values of each admittance component. [koziel2013surrogate]. This lumped circuit model, initially used in D-RIS to capture the relationship between reflection amplitude and phase shift for each reflecting element [practicalRIS_Abeywickrama], is generalized in BD-RIS to model the losses of each admittance component. The admittance of this equivalent circuit is given by
$Y_{m_{g},n_{g}}(C_{m_{g},n_{g}})=\frac{1}{\jmath\omega L_{1}}+\frac{1}{\jmath\omega L_{2}+\frac{1}{\jmath\omega C_{m_{g},n_{g}}}+R},~\forall m_{g},n_{g},$ (7)
@@
$\left(\Re\{Y_{m_{g},n_{g}}\}-\frac{1}{2R}\right)^{2}$ (9)
$\quad+\left(\Im\{Y_{m_{g},n_{g}}\}-\left(-\frac{1}{\omega L_{1}}\right)\right)^{2}=\left(\frac{1}{2R}\right)^{2}.$
To the best of our knowledge, the circular relationship in (9) has not been reported in prior literature and is analytically derived in this paper for the first time. Clearly, the radius of this circle is $\frac{1}{2R}$, which depends solely on the resistance $R$, and the center of the circle is located at $\left(\frac{1}{2R},-\frac{1}{\omega L_{1}}\right)$, which depends jointly on the resistance $R$ and the inductance $L_{1}$. In this paper, we utilize the parameters from the data sheet of the practical varactor diode SMV2020-079LF to model the equivalent circuit of each admittance component. Specifically, we set $L_{2}=0.7$ nH, $C_{m_{g},n_{g}}\in[0.35,3.20]$ pF [^4] The limited range of capacitance values, along with the presence of $L_{1}$ prevents the tunable admittances of interconnections from being completely ”turned off”, indicating that single-connected architecture is not a subset of the fully-/group-connected and tree-/forest-connected architectures in BD-RIS. and assume $L_{1}=6$ nH [^5] In [practicalRIS_Abeywickrama], $L_{1}$ is assumed to be 2.5 nH to achieve the full reflection tuning range for each reflecting element. However, this assumption is unnecessary in our BD-RIS model due to the nonlinear mapping from the admittance matrix to scattering matrix. The selection of the $L_{1}$ value is discussed in Section V.. As such, we represent the possible values of $Y_{m_{g},n_{g}}$ through the set $\mathcal{Y}$, defined as
$\mathcal{Y}=\{\alpha+\beta e^{\jmath\theta}\ \mid\ \theta\in[\theta_{\mathrm{min}},\theta_{\mathrm{max}}]\},$ (10)
@@
Remark 1: In most existing D-RIS [smart_radio_Marco_Di_renzo, IRS_magazine_Qingqing_Wu, IRS_tutorial_Qingqing_Wu, cascaded_CE_ris_Guo, weighted_sum_rate_ris_Guo] and most existing BD-RIS literature666[lossy_interconnections_nerini] studies lossy interconnections between admittance components while all the admittance components themselves are assumed to be lossless. The optimization methods proposed in [lossy_interconnections_nerini] are not applicable to our work, as we focus on BD-RIS with lossy tunable admittance components. [shen2021modelling, BD-RIS_magazine, discrete_nerini, close_form_nerini, hybrid_hongyu, multisector_hongyu, nerinigraphtheory, channel_estimation_hongyu, non_reciprocal_hongyu, universal_framework_nerini, mutual_coupling_hongyu, mutual_coupling_nerini, wideband_circuit_hongyu, universal_optimization_zheyu, dynamic_grouping_hongyu, band_stem_zheyu, parteo_frontier_nerini], the reconfigurable admittance network is assumed to be lossless, such that each tunable admittance is purely imaginary with $\Re\{Y_{m_{g},n_{g}}\}=0$ and the feasible range of its imaginary part lies unconstrained along the y-axis as illustrated in Fig. 2. This simple assumption fails to capture the coupling between the real and imaginary parts of $Y_{m_{g},n_{g}}$, which is important in practical realizations of BD-RIS and will have a non-negligible impact on the system performance as it will be illustrated in the following sections.
@@ III-A System Model and Problem Formulation
We start by considering an $M$-element BD-RIS-aided SISO wireless communication system. In this paper, we adopt the common assumptions from related literature [shen2021modelling, BD-RIS_magazine, discrete_nerini, close_form_nerini, hybrid_hongyu, multisector_hongyu, nerinigraphtheory, channel_estimation_hongyu, lossy_interconnections_nerini, wideband_circuit_hongyu, universal_optimization_zheyu, dynamic_grouping_hongyu, band_stem_zheyu, parteo_frontier_nerini], which state that the antennas at the transmitter, RIS, and receiver are perfectly matched, have no mutual coupling and no structural scattering. The overall channel $h$ between the transmitter and receiver can thus be written as
@@
Next, we aim to maximize the received signal power for the considered BD-RIS-aided SISO system, taking into account the losses in BD-RIS. As discussed in Section II-A, it is convenient to capture the losses of BD-RIS with admittance parameter. We introduce the following notation to define the BD-RIS architecture:
[…]
$\text{ are not connected}\big\}.$
In particular, $\mathcal{I}_{g}=\emptyset$ for group-connected BD-RIS, $\mathcal{I}_{g}=\{(m,n)\mid|m-n|>1\}$ for tridiagonal forest-connected BD-RIS, and $\mathcal{I}_{g}=\{(m,n)\mid m,n\neq c_{g},m\neq n\}$ for arrowhead forest-connected BD-RIS. By employing the relationship between the scattering matrix $\bm{\Phi}$ and the admittance matrix $\mathbf{Y}$ in (1), the mapping between $\mathbf{Y}$ and the tunable admittances $Y_{m_{g},n_{g}},~\forall m_{g},n_{g}$ in (2) – (4), and the set of feasible values for each $Y_{m_{g},n_{g}}$ in (10), the power maximization problem can be formulated as follows [^7] We include $\bm{\Phi},\{\mathbf{Y}_{g}\}$ and $\{Y_{m_{g},n_{g}}\}$ as optimization variables because they are treated as independent variables in the optimization process.:
@@
$\mathbf{Y}_{g}=\mathbf{Y}_{g}^{T},~\forall g\in\mathcal{G},$ (12d)
@@
$Y_{m_{g},n_{g}}=0,~\forall(m_{g},n_{g})\in\mathcal{I}_{g},$
$Y_{m_{g},n_{g}}\in\mathcal{Y},~\forall(m_{g},n_{g})\notin\mathcal{I}_{g}.$ (12f)
Problem (12) is challenging to solve because constraint (12b) introduces a nonlinear mapping between the admittance matrix and the scattering matrix, resulting from matrix inversion. Additionally, constraint (12f) is non-convex, which further complicates the optimization process. In the following subsections, we propose two solutions to maximize the received signal power with a lossy BD-RIS in the SISO systems. First, we propose an $M$-ADMM algorithm in Section III-B to directly solve (12). In Section III-C, we further propose a low-complexity algorithm that approximately maximizes the received signal power without directly solving (12).
Remark 2: In lossless BD-RIS modeling, constraint (12e) is typically not considered because there are no restrictions on individual tunable components $Y_{m_{g},n_{g}},~\forall m_{g},n_{g}$. Furthermore, constraint (12f) arises specifically from modeling lossy tunable admittance components, which does not exist in lossless BD-RIS models. More specifically, take a lossless fully-connected BD-RIS as an example, each $Y_{m_{g},n_{g}}$ has zero real part and arbitrary imaginary part, which leads to $\mathbf{Y}=\jmath\mathbf{B}$, with $\mathbf{B}$ being an arbitrary real-valued matrix. Consequently, the scattering matrix $\bm{\Phi}$ is an arbitrary unitary matrix. Therefore, one can focus on the design of either unconstrained $\mathbf{B}$ or constrained $\bm{\Phi}$ for lossless BD-RIS. However, neither of the above two strategies work for lossy BD-RIS since the real and imaginary parts of $Y_{m_{g},n_{g}}$ is inherently coupled and constrained by (12f). These two constraints (12e) and (12f), together with the nonlinear relation in (12b), constitute the main challenges introduced when accounting for lossy tunable components compared to lossless BD-RIS models.
III-B Proposed $M$-ADMM Algorithm
@@
Let $\bar{\mathbf{Y}}_{g}\in\mathbb{C}^{\bar{M}\times\bar{M}}$ be a matrix with entries $\bar{\mathbf{Y}}_{g}]_{m,n}=Y_{m_{g},n_{g}}$, then (12e) can be expressed as a linear mapping from $\bar{\mathbf{Y}}_{g}$ to $\mathbf{Y}_{g},~\forall g\in\mathcal{G}$. To facilitate the optimization, we further re-write this linear mapping into a vector form as $\mathsf{vec}(\mathbf{Y}_{g})=\mathbf{B}\mathsf{vec}(\bar{\mathbf{Y}}_{g})$, where matrix $\mathbf{B}\in\{0,\pm 1\}^{\bar{M}^{2}\times\bar{M}^{2}}$ maps $\mathsf{vec}(\bar{\mathbf{Y}}_{g})$ to $\mathsf{vec}(\mathbf{Y}_{g})$, $\forall g\in\mathcal{G}$, and is defined as
@@
$\forall i,j,k\in\mathcal$$\bar${M}}. It is worth noting that $\bar{\mathbf{Y}}_{g},~\forall g\in\mathcal{G}$, is symmetric due to the symmetry of $\mathbf{Y}_{g}$ and the relation between $\mathbf{Y}_{g}$ and $\bar{\mathbf{Y}}_{g}$ in (12e). In addition, the BD-RIS architecture determines the locations of non-zero elements in $\bar{\mathbf{Y}}_{g}$; see (12f). Hence, the actual degrees of freedom to be designed in $\bar{\mathbf{Y}}_{g}$ are the non-zero entries in its upper-triangular part, whose cardinality equals the number of tunable admittance components in each BD-RIS group, denoted by $U$. Let $\bar{\mathbf{y}}_{g}\in\mathbb{C}^{U\times 1}$ be the vector collecting these entries. The symmetry and architecture-related constraints can then be jointly incorporated by rewriting $\mathsf{vec}(\bar{\mathbf{Y}}_{g})=\mathbf{P}\bar{\mathbf{y}}_{g},~\forall g\in\mathcal{G}$, where $\mathbf{P}\in\{0,1\}^{\bar{M}^{2}\times U}$ is a binary matrix mapping $\bar{\mathbf{y}}_{g}$ to $\mathsf{vec}(\bar{\mathbf{Y}}_{g})$. We provide the explicit forms of matrix $\mathbf{P}$ and $U$ for the group-connected and forest-connected architectures in the Appendix.
According to the aforementioned discussions, problem (15) can be equivalently re-written as
@@
$\mathrm{s.t.}$~~~~~~ $\bm{\Phi}_{g}=\left(Y_{0}\mathbf{I}_{\bar{M}}+\mathbf{Y}_{g}\right)^{-1}\left(Y_{0}\mathbf{I}_{\bar{M}}-\mathbf{Y}_{g}\right),~\forall g\in\mathcal{G},$ (17b)
$\mathbf{Y}_{g}=\overline{\mathsf{vec}}(\mathbf{B}\mathbf{P}\bar{\mathbf{y}}_{g}),~\forall g\in\mathcal{G},$ (17c)
$[\bar{\mathbf{y}}_{g}]_{m^{\prime}}\in\mathcal{Y},~\forall g\in\mathcal{G},~\forall m^{\prime}\in\mathcal{M}^{\prime},$ (17d)
where $\mathcal{M^{\prime}}=\{1,\ldots,U\}$. To address this, we propose a double-loop iterative algorithm, referred to as the $M$-ADMM algorithm. In the outer loop, we utilize the $M$ method [$M$] to linearize objective (17a). Then, in the inner loop, we employ the ADMM method [boyd_admm] to effectively handle the complex constraints. The details of the proposed algorithm are described as follows.
III-B.1 $M$ Approximation
The main idea of the $M$ method for solving (17) is to replace the objective function with a locally tight lower-bound surrogate function, usually a concave one, such that the resulting problem is easier to tackle. For (17a), a simple and common approach for this approximation is to use first-order Taylor expansion of $\tilde{\bm{\phi}}^{H}\mathbf{F}\tilde{\bm{\phi}}$. Defining $g(\tilde{\bm{\phi}})=\tilde{\bm{\phi}}^{H}\mathbf{F}\tilde{\bm{\phi}}$, at iteration $t$, we have